OpenStudy (anonymous):

what is the solution to the equation x-10=6

OpenStudy (anonymous):

Thats math hunny..

OpenStudy (anonymous):

hahaha smile

OpenStudy (anonymous):

btw the answer is 16... you have to isolate x by getting rid of the surrounding numbers... so you would add 10 to both sides to get rid of it. <- idk if that makes sense

OpenStudy (anonymous):

x-10=6 +10=+10 x=16